# Restructuring Tractable Probabilistic Circuits

### v1

@@ 3.3 Computing Vtree Labelling
The next question that immediately arises is how to compute a valid labelling for $W$ with respect to $G_{\mathscr{A}}$. One naive solution is to set $\mathbf{C}_{w}$ to be $\mathbf{Z}$, the set of all latent variables in $G_{\mathscr{A}}$. However, this is not desirable as $M^{\prime}\!=\!\max_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|\!=\!|\mathbf{Z}|\!=\!n\!-\!1$, resulting in the restructured circuit having exponential size $O(nh^{n-1})$. Hence we present a greedy approach that computes a labelling while trying to minimize $M^{\prime}$.
The algorithm proceeds top-down on $W$. For the base case where $w$ is the root, we set $\mathbf{C}_{w}:=\varnothing$. For the inductive step, let $l$ and $r$ be the children of $w$ and assume that we have computed $\mathbf{C}_{w}$ as a cover for $\mathbf{X}_{w}$ in $G_{\mathscr{A}}$: we (1) split $G_{\mathscr{A}}$ into connected components $\{G_{i}\}$ via $\mathbf{C}_{w}$; then (2) within each connected component $G_{i}$, we compute a minimum d-separator $\mathbf{C}_{i}$ that blocks all paths between $\mathbf{X}_{l}\cap G_{i}$ and $\mathbf{X}_{r}\cap G_{i}$ by calling the sub-routine MinimumSeparator. We set $\mathbf{D}_{w}:=\left({\bigcup}_{i}\mathbf{C}_{i}\right)\cup\mathbf{C}_{w}$ and observe that $\mathbf{D}_{w}$ covers both $\mathbf{X}_{l}$ and $\mathbf{X}_{r}$ in $G_{\mathscr{A}}$. To compute $\mathbf{C}_{l}$, similarly for $\mathbf{C}_{r}$, we consider all paths starting from $\mathbf{X}_{l}$ and stopping immediately when reaching some $Z_{j}\in\mathbf{D}_{w}$, and we let $\mathbf{C}_{l}$ to be the set containing all such $Z_{j}$s. The pseudo code is shown in Algorithm 2.
Algorithm 2 Computing $\mathbf{C}_{w}$ for $w\!\in\!W$
procedure ComputeLabel($w,\mathbf{C}_{w}$)
     $\{G_{i}\}\leftarrow\textsc{ConnectedComponents}(G_{\mathscr{A}},\mathbf{C}_{w})$
     $\mathbf{C}_{i}\leftarrow\textsc{MinimumSeparator}(G_{i},\mathbf{X}_{l}\!\cap\!G_{i},\mathbf{X}_{r}\!\cap\!G_{i})$
     $\mathbf{D}_{w}\leftarrow\left({\bigcup}_{i}\mathbf{C}_{i}\right)\cup\mathbf{C}_{w}$
     $\mathbf{C}_{l}\leftarrow\{Z_{j}\!\in\!\mathbf{D}_{w}:\textsc{Paths}(\mathbf{X}_{l},Z_{j})\cap\mathbf{D}_{w}\!=\!\{Z_{j}\}\}$
     $\mathbf{C}_{r}\leftarrow\{Z_{j}\!\in\!\mathbf{D}_{w}:\textsc{Paths}(\mathbf{X}_{r},Z_{j})\cap\mathbf{D}_{w}\!=\!\{Z_{j}\}\}$
     $\textsc{ComputeLabel}(l,\mathbf{C}_{l})$
     $\textsc{ComputeLabel}(r,\mathbf{C}_{r})$
@@
Note that the MinimumSeparator procedure called in Algorithm 2 computes a minimum d-separator that blocks all paths between $\mathbf{X}_{l}$ and $\mathbf{X}_{r}$ in the subgraph $G_{i}$. Even though polytime algorithms for computing minimum d-separators exist in literature (Tian et al.,, 1998), we derive a linear-time algorithm that is easy to implement for our use case, where $G_{i}$ is a rooted tree with leaves in $\mathbf{X}_{l}$, $\mathbf{X}_{r}$ and $\mathbf{C}_{w}$. We refer readers to the Appendix for details.
@@
We prove by a top-down induction on $W$ that the labelling $\mathbf{C}_{w}$ computed by Algorithm 2 is valid. Assume that $\mathbf{C}_{w}$ covers $\mathbf{X}_{w}$ in $G_{\mathscr{A}}$, we want to show that $\mathbf{C}_{l}$ and $\mathbf{C}_{r}$ satisfy the properties from Definition 3.6. To prove that $\mathbf{C}_{l}$ covers $\mathbf{X}_{l}$, we consider a path from $X_{a}\!\in\!\mathbf{X}_{l}$ to $X_{b}\!\in\!\mathbf{X}\setminus\mathbf{X}_{l}$. (1) If $X_{a}$ and $X_{b}$ are in the same $G_{i}$, then the path is blocked by $\mathbf{C}_{i}$. (2) If $X_{a}$ and $X_{b}$ are in different $G_{i}$s, then the path contains some node $Z\in\mathbf{C}_{w}$, and we can choose from the path the first $Z\!\in\!\mathbf{C}_{w}$. Then $Z\!\in\!\mathbf{C}_{l}$ by construction, implying that the path is blocked by $\mathbf{C}_{l}$. Hence we conclude that $\mathbf{C}_{l}$ is a cover for $\mathbf{X}_{l}$, satisfying Property 1. To prove that $\mathbf{C}_{l}$ satisfies Property 2, we argue that because $\mathbf{C}_{r}$ and $\mathbf{C}_{w}$ are both subsets of $\mathbf{D}_{w}$, all paths from $\mathbf{X}_{l}$ to $\mathbf{C}_{r}\cup\mathbf{C}_{w}$ will be blocked by $\mathbf{C}_{l}$ by the way that $\mathbf{C}_{l}$ is constructed. We can show that $\mathbf{C}_{r}$ satisfies Property 1 and Property 3 by the same argument. ∎
Though Algorithm 2 computes a valid labelling while greedily minimizing $|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|$, we do not know whether $M^{\prime}=\max_{w\in W}|\mathbf{C}_{l}\cup\mathbf{C}_{r}\cup\mathbf{C}_{w}|$ is globally minimized or not. In addition, we hypothesize that if we can find a minimum vtree labelling, then the size of the PC constructed by Algorithm 1 is optimal. We leave it as an open problem to design an algorithm that computes minimum labellings and prove the optimality of Algorithm 1 given a minimum labelling.
@@ 3.4 Corollaries
A sum node is deterministic if for every value $\mathbf{x}$ of $\mathbf{X}$, at most one child $c$ returns a non-zero value (i.e. $p_{c}(\mathbf{x})>0$). A PC is determinstic if all of its sum nodes are deterministic.
@@ 4 PC MULTIPLICATION
One important application of restructuring PCs is circuit multiplication: given two PCs $\mathscr{A}$ and $\mathscr{B}$, the goal is to construct a tractable PC $\mathscr{C}$ such that $p_{\mathscr{C}}(\mathbf{x})\propto p_{\mathscr{A}}(\mathbf{x})\cdot p_{\mathscr{B}}(\mathbf{x})$. PC multiplication was previously only addressed for structured PCs respecting the same vtree (Shen et al.,, 2016; Vergari et al.,, 2021). Circuit restructuring immediately gives us a means of multiplying two structured circuits respecting different vtrees, as we can simply restructure one of them to be compatible with the other. Though the restructured PC will in general have exponential size, in this section, we consider practical cases where circuit multiplications with respect to different vtrees is tractable.
We start by introducing a new structural property of tractable PCs called contiguity.
@@
Note that a contiguous circuit is not necessarily structured-decomposable and $0.5\otimes p(X_{1})\otimes p(X_{2},X_{3})\oplus 0.5\otimes p(X_{1},X_{2})\otimes p(X_{3})$ is such an example. Intuitively, random variables forming contiguous scopes can often be covered by vtree labellings of small cardinalities.
@@
Case 1. For the multiplication of contiguous PCs $\mathscr{A}$ and $\mathscr{B}$, we start by considering the case when $\mathscr{A}$ is a contiguous structured PC respecting the linear vtree $V$ and $\mathscr{B}$ is a contiguous structured PC respecting an arbitrary vtree $W$. It follows from Section 3.1 that the Bayesian network representation for $\mathscr{A}$ is a hidden Markov model (Rabiner,, 1989), as shown in Figure 3. By the definition of contiguity, each node $w\!\in\!W$ has a scope of the form $\mathbf{X}_{a:b}:=\{X_{a},\dots,X_{b}\}$ and we can label it with $\mathbf{C}_{a:b}:=\{Z_{a},Z_{b+1}\}$; in particular, we drop $Z_{a}$ if $a=1$ and drop $Z_{b+1}$ if $b=n$.
@@
$\mathbf{C}_{a:b}$ is a valid vtree labelling of $W$ respecting $G_{\mathscr{A}}$ with cardinality $M^{\prime}=3$.
@@
Case 2. Then we consider the more general case where $\mathscr{A}$ is a contiguous structured PC of depth $d$ respecting vtree $V$ and $\mathscr{B}$ is a contiguous structured PC with an arbitrary vtree $W$. Similarly to the previous case, our goal is to come up with a small labelling of $W$ with respect to $G_{\mathscr{A}}$. Since $\mathscr{A}$ is contiguous, its vtree $V$ can be viewed as a segment tree (Cormen et al.,, 2022). Algorithm 3, which is adapted from the segment tree querying algorithm, computes a cover $\mathbf{C}_{a:b}\subseteq G_{\mathscr{A}}$ for each contiguous segment $\mathbf{X}_{a:b}$. For each $w\!\in\!W$, $\mathbf{X}_{w}=X_{a:b}$ for some $1\leq a\leq b\leq n$ and we set $\mathbf{C}_{w}=\mathbf{C}_{a:b}=\textsc{SegmentCover}(V,X_{a:b})$.
Algorithm 3 Compute Cover for Segment $\mathbf{X}_{a:b}$
procedure SegmentCover($v$, $\mathbf{X}_{a:b}$)
     if $\mathbf{X}_{a:b}=\varnothing$ then
@@
     $\mathbf{L}\leftarrow\textsc{SegmentCover}(l,\mathbf{X}_{l}\cap\mathbf{X}_{a:b})$
     $\mathbf{R}\leftarrow\textsc{SegmentCover}(r,\mathbf{X}_{r}\cap\mathbf{X}_{a:b})$
     return $\mathbf{L}\cup\mathbf{R}$
@@
$\mathbf{C}_{w}$ is a valid vtree labelling with respect to $G_{\mathscr{A}}$.
In addition to the fact that $\mathbf{C}_{w}$ is a valid labelling, by the runtime analysis of the original segment tree querying algorithm, we know that the number of nodes visited at each level of $V$ is at most $4$ and it follows that $|\mathbf{C}_{w}|\leq 4d$; hence the cardinality of $\mathbf{C}_{w}$ is bounded by $12d$. Then following a similar analysis, we have the following results for multiplying two contiguous PCs.
@@
Remark 4.8.
Thus far, we have assumed that both $\mathscr{A}$ and $\mathscr{B}$ are structured PCs. We claim that we can further generalize our results by removing the constraint that $\mathscr{B}$ has to be structured, and Theorems 4.3 and 4.5 would still hold. We illustrate the idea by showing how to multiply a contiguous structured PC $\mathscr{A}$ respecting a linear vtree and an arbitrary contiguous PC $\mathscr{B}$. Since $\mathscr{B}$ is not structured decomposable, we cannot restructure $\mathscr{A}$ to the vtree of $\mathscr{B}$. However, we can use the same idea as Algorithm 1 to restructure $\mathscr{A}$ “on-the-fly” as we multiply it with $\mathscr{B}$ in a bottom-up way. Specifically, for each possible scope $\mathbf{X}_{a:b}$ that appears in $\mathscr{B}$, we recursively construct circuit representations for the functions $p_{q}(\mathbf{X}_{a:b})\cdot p_{\mathscr{A}}(\mathbf{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ for $i,j$ and $q\!\in\!\mathscr{B}$ with scope $\mathbf{X}_{a:b}$. The recurrence relation is similar to that of Algorithm 1 and we refer readers to the Appendix for details.
As an explicit application of circuit multiplication, let us consider constrained text generation (Zhang et al.,, 2024), in which linear PCs (HMMs) are multiplied with deterministic finite automata (DFAs) representing the logical constraint. Our results imply that one can also multiply a HMM with a contiguous logical circuit such as a sentential decision diagram (SDD) (Darwiche,, 2011), which have been shown to be exponentially more expressive efficient (Bova,, 2016).

### v2

@@ 3.3 Computing Vtree Labelling
The next question that immediately arises is how to compute a valid labelling for $W$ with respect to $G_{\mathscr{A}}$. One naive solution is to set $\bm{C}_{w}$ to be $\bm{Z}$, the set of all latent variables in $G_{\mathscr{A}}$. However, this is not desirable as $M^{\prime}\!=\!\max_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|\!=\!|\bm{Z}|\!=\!n\!-\!1$, resulting in the restructured circuit having exponential size $O(nh^{n-1})$. Hence we present a greedy approach that computes a labelling while trying to minimize $M^{\prime}$.
The algorithm proceeds top-down on $W$. For the base case where $w$ is the root, we set $\bm{C}_{w}:=\varnothing$. For the inductive step, let $l$ and $r$ be the children of $w$ and assume that we have computed $\bm{C}_{w}$ as a cover for $\bm{X}_{w}$ in $G_{\mathscr{A}}$: we (1) split $G_{\mathscr{A}}$ into connected components $\{G_{i}\}$ via $\bm{C}_{w}$; then (2) within each connected component $G_{i}$, we compute a minimum d-separator $\bm{C}_{i}$ that blocks all paths between $\bm{X}_{l}\cap G_{i}$ and $\bm{X}_{r}\cap G_{i}$ by calling the sub-routine MinimumSeparator. We set $\bm{D}_{w}:=\left({\bigcup}_{i}\bm{C}_{i}\right)\cup\bm{C}_{w}$ and observe that $\bm{D}_{w}$ covers both $\bm{X}_{l}$ and $\bm{X}_{r}$ in $G_{\mathscr{A}}$. To compute $\bm{C}_{l}$, similarly for $\bm{C}_{r}$, we consider all paths starting from $\bm{X}_{l}$ and stopping immediately when reaching some $Z_{j}\in\bm{D}_{w}$, and we let $\bm{C}_{l}$ to be the set containing all such $Z_{j}$s. The pseudo code is shown in Algorithm 2.
Algorithm 2 Computing $\bm{C}_{w}$ for $w\!\in\!W$
procedure ComputeLabel($w,\bm{C}_{w}$)
     $\{G_{i}\}\leftarrow\textsc{ConnectedComponents}(G_{\mathscr{A}},\bm{C}_{w})$
     $\bm{C}_{i}\leftarrow\textsc{MinimumSeparator}(G_{i},\bm{X}_{l}\!\cap\!G_{i},\bm{X}_{r}\!\cap\!G_{i})$
     $\bm{D}_{w}\leftarrow\left({\bigcup}_{i}\bm{C}_{i}\right)\cup\bm{C}_{w}$
     $\bm{C}_{l}\leftarrow\{Z_{j}\!\in\!\bm{D}_{w}:\textsc{Paths}(\bm{X}_{l},Z_{j})\cap\bm{D}_{w}\!=\!\{Z_{j}\}\}$
     $\bm{C}_{r}\leftarrow\{Z_{j}\!\in\!\bm{D}_{w}:\textsc{Paths}(\bm{X}_{r},Z_{j})\cap\bm{D}_{w}\!=\!\{Z_{j}\}\}$
     $\textsc{ComputeLabel}(l,\bm{C}_{l})$
     $\textsc{ComputeLabel}(r,\bm{C}_{r})$
@@
Note that the MinimumSeparator procedure called in Algorithm 2 computes a minimum d-separator that blocks all paths between $\bm{X}_{l}$ and $\bm{X}_{r}$ in the subgraph $G_{i}$. Even though polytime algorithms for computing minimum d-separators exist in literature (Tian et al.,, 1998), we derive a linear-time algorithm that is easy to implement for our use case, where $G_{i}$ is a rooted tree with leaves in $\bm{X}_{l}$, $\bm{X}_{r}$ and $\bm{C}_{w}$. We refer readers to the Appendix for details.
@@
We prove by a top-down induction on $W$ that the labelling $\bm{C}_{w}$ computed by Algorithm 2 is valid. Assume that $\bm{C}_{w}$ covers $\bm{X}_{w}$ in $G_{\mathscr{A}}$, we want to show that $\bm{C}_{l}$ and $\bm{C}_{r}$ satisfy the properties from Definition 3.7. To prove that $\bm{C}_{l}$ covers $\bm{X}_{l}$, we consider a path from $X_{a}\!\in\!\bm{X}_{l}$ to $X_{b}\!\in\!\bm{X}\setminus\bm{X}_{l}$. (1) If $X_{a}$ and $X_{b}$ are in the same $G_{i}$, then the path is blocked by $\bm{C}_{i}$. (2) If $X_{a}$ and $X_{b}$ are in different $G_{i}$s, then the path contains some node $Z\in\bm{C}_{w}$, and we can choose from the path the first $Z\!\in\!\bm{C}_{w}$. Then $Z\!\in\!\bm{C}_{l}$ by construction, implying that the path is blocked by $\bm{C}_{l}$. Hence we conclude that $\bm{C}_{l}$ is a cover for $\bm{X}_{l}$, satisfying Property 1. To prove that $\bm{C}_{l}$ satisfies Property 2, we argue that because $\bm{C}_{r}$ and $\bm{C}_{w}$ are both subsets of $\bm{D}_{w}$, all paths from $\bm{X}_{l}$ to $\bm{C}_{r}\cup\bm{C}_{w}$ will be blocked by $\bm{C}_{l}$ by the way that $\bm{C}_{l}$ is constructed. We can show that $\bm{C}_{r}$ satisfies Property 1 and 3 by the same argument. ∎
As an example, we (partially) illustrate the application of Algorithm 2 to the target vtree in Figure 1(c). For node $a$ and its children $b,c$, the only connected component is $G_{1}:=G_{\mathscr{A}}$. A minimum separator of $\bm{X}_{b}=\{X_{3},X_{6},X_{4},X_{5}\}$ and $\bm{X}_{c}=\{X_{1},X_{2}\}$ is then given by $\bm{C}_{1}:=\{Z_{1:3},Z_{2:3}\}$. Then $\bm{D}_{1}=\bm{C}_{1}\cup\bm{C}_{a}=\{Z_{1:3},Z_{2:3}\}$ covers both $\bm{X}_{b},\bm{X}_{c}$. Tracing the paths from each $X$ variable to $\bm{D}_{1}$ finally gives $\bm{C}_{b}=\bm{C}_{c}=\{Z_{1:3},Z_{2:3}\}$.
Though Algorithm 2 computes a valid labelling while greedily minimizing $|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|$, we do not know whether $M^{\prime}=\max_{w\in W}|\bm{C}_{l}\cup\bm{C}_{r}\cup\bm{C}_{w}|$ is globally minimized or not. In addition, we hypothesize that if we can find a minimum vtree labelling, then the size of the PC constructed by Algorithm 1 is optimal. We leave it as an open problem to design an algorithm that computes minimum labellings and prove the optimality of Algorithm 1 given a minimum labelling.
@@ 3.4 Corollaries
A sum node is deterministic if for every value $\bm{x}$ of $\bm{X}$, at most one child $c$ returns a non-zero value (i.e. $p_{c}(\bm{x})>0$). A PC is determinstic if all of its sum nodes are deterministic.
@@
Contiguous & StructuredLinearLog-Depth & ContiguousAny Log-DepthContiguousPolynomialQuasi-Polynomial
Figure 3: Summary of restructuring results; the top/bottom rows indicate the source/target structures. Blue and bold arrows indicate novel complexity results. For the two other arrows that were known to be polynomial-time, our approach yields more efficient algorithms amenable to practical implementation.
@@ 4 PC MULTIPLICATION
One important application of restructuring PCs is circuit multiplication: given two PCs $\mathscr{A}$ and $\mathscr{B}$, the goal is to construct a tractable PC $\mathscr{C}$ such that $p_{\mathscr{C}}(\bm{x})\propto p_{\mathscr{A}}(\bm{x})\cdot p_{\mathscr{B}}(\bm{x})$. PC multiplication was previously only addressed for structured PCs respecting the same vtree (Shen et al.,, 2016; Vergari et al.,, 2021). Circuit restructuring immediately gives us a means of multiplying two structured circuits respecting different vtrees, as we can simply restructure one of them to be compatible with the other. Though the restructured PC will in general have exponential size, in this section, we consider practical cases where circuit multiplications with respect to different vtrees is tractable. Further, we will show how “on-the-fly” restructuring can enable circuit multiplication even when one of the circuits is not structured-decomposable.
We start by introducing a new structural property of tractable PCs called contiguity.444 Amarilli et al., (2017) defined a similar concept called compatible order in the context of logical circuits.
@@
Contiguous PCs subsume many widely-used subclasses of PCs such as Hidden Markov Models (HMMs). Note that a contiguous circuit is not necessarily structured-decomposable and $0.5\otimes p(X_{1})\otimes p(X_{2},X_{3})\oplus 0.5\otimes p(X_{1},X_{2})\otimes p(X_{3})$ is such an example. The fact that contiguous PCs need not be structured-decomposable also suggests new designs of circuits that can be tractably multiplied. One important example is probabilistic context-free grammars (PCFGs).
Theorem 4.2.
Let $G$ be a PCFG consisting of $m$ production rules. For sequences of length $n$, $G$ can be represented as a contiguous yet non-structured PC $\mathscr{A}$ of size $O(mn^{3})$; i.e. $p_{\mathscr{A}}(\bm{x})$ computes the probability of $\bm{x}$ being derived from $G$, for all $\bm{x}$ of length $n$.
The algorithm for constructing the PC representation for a PCFG is in spirit similar to the CYK algorithm (Kasami,, 1966; Younger,, 1967; Cocke,, 1969): given a PCFG, for each contiguous subsegment $[a,b]$ of $1,2,\dots,n$, and each non-terminal $N$ in the PCFG, we can construct a sum node $u$ (recursively) such that the sub-circuit rooted at $u$ represents the distribution over the set of strings on the segment $[a,b]$ that can be derived from $N$.
In the following, we consider the problem of multiplying two contiguous PCs $\mathscr{A}$ and $\mathscr{B}$. Intuitively, random variables forming contiguous scopes are more likely to be covered by vtree labellings of small cardinalities, which would give efficient restructuring by Section 3. We start by considering the simpler case where both circuits are structured. We then generalize our results by allowing one of them to be non-structured. Figure 3 summarizes our main results (Theorems 4.4, 4.6, 4.9).
4.1 Multiplying Contiguous Structured PCs
@@
Case 1. For the multiplication of contiguous PCs $\mathscr{A}$ and $\mathscr{B}$, we start by considering the case when $\mathscr{A}$ is a contiguous structured PC respecting the linear vtree $V$ and $\mathscr{B}$ is a contiguous structured PC respecting an arbitrary vtree $W$. It follows from Section 3.1 that the Bayesian network representation for $\mathscr{A}$ is a hidden Markov model (Rabiner,, 1989), as shown in Figure 4. By the definition of contiguity, each node $w\!\in\!W$ has a scope of the form $\bm{X}_{a:b}:=\{X_{a},\dots,X_{b}\}$ and we can label it with $\bm{C}_{a:b}:=\{Z_{a},Z_{b+1}\}$; in particular, we drop $Z_{a}$ if $a=1$ and drop $Z_{b+1}$ if $b=n$.
@@
$\bm{C}_{a:b}$ is a valid vtree labelling of $W$ respecting $G_{\mathscr{A}}$ with cardinality $M^{\prime}=3$.
@@
Case 2. Then we consider the more general case where $\mathscr{A}$ is a contiguous structured PC of depth $d$ respecting vtree $V$ and $\mathscr{B}$ is a contiguous structured PC with an arbitrary vtree $W$. Similarly to the previous case, our goal is to come up with a small labelling of $W$ with respect to $G_{\mathscr{A}}$. Since $\mathscr{A}$ is contiguous, its vtree $V$ can be viewed as a segment tree (Cormen et al.,, 2022). Algorithm 3, which is adapted from the segment tree querying algorithm, computes a cover $\bm{C}_{a:b}\subseteq G_{\mathscr{A}}$ for each contiguous segment $\bm{X}_{a:b}$. For each $w\!\in\!W$, $\bm{X}_{w}=X_{a:b}$ for some $1\leq a\leq b\leq n$ and we set $\bm{C}_{w}=\bm{C}_{a:b}=\textsc{SegmentCover}(V,X_{a:b})$.
Algorithm 3 Compute Cover for Segment $\bm{X}_{a:b}$
procedure SegmentCover($v$, $\bm{X}_{a:b}$)
     if $\bm{X}_{a:b}=\varnothing$ then
@@
     $\bm{L}\leftarrow\textsc{SegmentCover}(l,\bm{X}_{l}\cap\bm{X}_{a:b})$
     $\bm{R}\leftarrow\textsc{SegmentCover}(r,\bm{X}_{r}\cap\bm{X}_{a:b})$
     return $\bm{L}\cup\bm{R}$
@@
$\bm{C}_{w}$ is a valid vtree labelling with respect to $G_{\mathscr{A}}$, and $|\bm{C}_{w}|\!\leq\!4d$ for all $w$.
In addition to the fact that $\bm{C}_{w}$ is a valid labelling, $|\bm{C}_{w}|\!\leq\!4d$ follows from the fact that the number of nodes visited at each level of $V$ is at most $4$ (see Appendix); hence the cardinality of $\bm{C}_{w}$ is bounded by $12d$. Then following a similar analysis, we have the following results for multiplying two contiguous PCs.
@@
4.2 Generalization to Non-structured PCs
Thus far, we have assumed that both $\mathscr{A}$ and $\mathscr{B}$ are structured PCs. We claim that we can further generalize our results by removing the constraint that $\mathscr{B}$ has to be structured, and Theorems 4.4 and 4.6 would still hold. We illustrate the idea by showing how to multiply a contiguous structured PC $\mathscr{A}$ respecting a linear vtree and an arbitrary contiguous PC $\mathscr{B}$. Since $\mathscr{B}$ is not structured decomposable, we cannot restructure $\mathscr{A}$ to the vtree of $\mathscr{B}$. However, we can use the same idea as Algorithm 1 to restructure $\mathscr{A}$ “on-the-fly” as we multiply it with $\mathscr{B}$ in a bottom-up way. Specifically, for each possible scope $\bm{X}_{a:b}$ that appears in $\mathscr{B}$, we recursively construct circuit representations for the functions $p_{q}(\bm{X}_{a:b})\cdot p_{\mathscr{A}}(\bm{X}_{a:b}\>|\>Z_{a}\!=\!i,Z_{b}\!=\!j)$ for $i,j$ and $q\!\in\!\mathscr{B}$ with scope $\bm{X}_{a:b}$. The recurrence relation is similar to that of Algorithm 1 and we refer readers to the Appendix for details.
Theorem 4.9.
As an explicit application of circuit multiplication, let us consider constrained text generation, in which linear PCs (HMMs) are multiplied with deterministic finite automata (DFAs) representing the logical constraint (Zhang et al.,, 2024). Our results imply that one can also multiply HMMs with contiguous logical circuits such as a sentential decision diagrams (SDDs) (Darwiche,, 2011), which have been shown to be exponentially more expressive efficient (Bova,, 2016). It also follows from Theorem 4.2 that we can tractably multiply HMMs with (P)CFGs, subsuming a recent result that HMMs can be tractably marginalized over unambiguous CFGs (Marzouk and de La Higuera,, 2022). Our results on multiplying contiguous PCs open up new possibilities for generalizing (Zhang et al.,, 2024) to much broader families of circuits and constraints.